# Crosscap Quenches and Entanglement Evolution

### v1

@@
Understanding the mechanisms by which complex correlations emerge through the dynamics of quantum many-body systems remains a fundamental challenge in modern physics. To address this, quench dynamics starting from nonequilibrium states have been extensively studied, leading to significant progress. In this paper, we propose a novel quench protocol, termed the “crosscap quench”, to investigate how highly structured thermal pure states relax into typical ones. We begin by analyzing conformal field theories (CFTs) and derive universal features in the time evolution of entanglement entropy. Furthermore, leveraging the AdS/CFT correspondence, we study holographic CFTs, providing an analytically tractable example in chaotic CFTs. Finally, we validate these findings through numerical simulations in both nonintegrable and integrable quantum spin systems.
@@ Introduction
While |$\ket{\rm EAP(\beta)}$ serves as a thermal pure state, it is a highly atypical one. In fact, it can be easily distinguished from a typical state by the antipodal correlation functions, which are defined on a few degrees of freedom but are considered highly non-local in 1D systems. In this sense, the EAP state utilizes the antipodal entanglement structure to deceive local observers into perceiving thermal equilibrium and at the same time stays easy to simulate.
As a result, even though |$\ket{\rm EAP(\beta)}$ is already in thermal equilibrium, its correlation structure will get more scrambled under chaotic quenches. We will study the time evolution of various type of correlations under such quenches, which will highlight a hierarchy between different notions of thermal equilibrium. We will also study time evolution of EAP states in integrable systems, see how the initial thermal condition deviates from the equilibrium.
@@
Note added: While this paper is under preparation, a very interesting paper [9] came out and also studied time evolution of entanglement entropy from crosscap states but in different systems. The authors of [9] focused on quantum circuit models and integrable fermion chains while we focus on CFT, holography and spin chains in this paper. The integrable fermion chains investigated in [9] should be equivalent to our computations for the transverse-field Ising models in section 5 via the Jordan-Wigner transformation in some regimes.
@@ 2.1 Entanglement Structure of Crosscap States
The first situation we would like to consider is a single interval subsystem $A$ given by $0\leq x\leq l$. At the thermodynamic limit limit $\alpha/L\ll 1$, in the regime $l/L\ll 1$,
@@ 3.4 Time Evolution of EE after the Crosscap Quench
Here, |$\ket{0}_{j}$ and |$\ket{1}_{j}$ are the eigenvectors of $\sigma_{j}^{z}$ in the local Hilbert space associated with site $j$. Consequently, the state at time $t$ is given by
@@
Figure 1 shows the dynamics of entanglement entropy for the single interval $A=\{1,2,\dots,l\}$ and the antipodally located double interval $A=\{1,2,\dots,l/2\}\cup\{N+1,N+2,\dots,N+l/2\}$. For $l\gg L$, the entanglement entropy exhibits an initial linear growth and saturates at a constant value, in well agreement with CFT predictions. Moreover, this saturated value coincides with the thermodynamic entropy at $\beta=0$. These results implies that the EAP state acquires a complex correlation structure and relaxes into a few-body thermal state.
Figure 1: Entanglement entropy $S_{A}$ of the time-evolved EAP state |$\ket{{\rm EAP}(t)}$ in the critical Heisenberg chain with next-nearest-neighbor interactions as a function of $t$ for total system size $202N=$ 20 and various subsystem size $l$, with solid lines representing the antipodally located double intervals $A=\{1,2,\dots,l/2\}\cup\{N+1,N+2,\dots,N+l/2\}$ and dashed lines represent the single intervals $A=\{1,2,\dots,l\}$.
As illustrated in Fig. 2, the time-evolved EAP state |$\ket{{\rm EAP}(t)}$ contains a volume-law entanglement, with a slight deviation increasing as the subsystem size grows, and the deviation becomes most pronounced when the subsystem size is exactly half of the total system size. Furthermore, we plot the system-size dependence of the deviation in Fig. 2. It can be seen that the deviation is subextensive. This is consistent with the results obtained from the analysis of holographic CFTs. Furthermore, we find that the deviation scales as $N^{1/2}$ in the thermodynamic limit.
Figure 2: (2) Time-averaged entanglement entropy of the time-evolved EAP state |$\ket{{\rm EAP}(t)}$ in the critical Heisenberg chain with next-nearest-neighbor interactions, averaged over the time interval $t\in[8,16]$, as a function of the subsystem size $l$ for the single interval $A$ in a system of size $202N=$ 20. The solid line represents the entanglement of the initial EAP state, which is maximally entangled for any single interval. (2) Deviation from the maximum value at $l=N$. The solid line serves as a guide to the eye and is proportional to $N^{-1/2}$.
@@
Figure 4: (4) Time-averaged entanglement entropy of the time-evolved EAP state |$\ket{{\rm EAP}(t)}$ in the critical transverse-field Ising chain, averaged over the time interval $t\in[8,32]$, as a function of the subsystem size $l$ for the single interval $A$ in a system of size $202N=$ 20. The solid line represents the entanglement of the initial EAP state, which is maximally entangled for any single interval. (4) Deviation from the maximum value at $l=N$.
@@ Conclusion
In this paper, we have studied quantum quenches in 1D many-body systems starting from crosscap states and EAP states. While EAP states are originally defined in spin systems, we have argued that crosscap states serve as good analogues of them in CFT because they exhibit similar entanglement structures. This was demonstrated by studying the universal behavior of the entanglement Rényi entropy of CFT crosscap states. Based on this, we termed the quenches starting from crosscap states and EAP states as ”crosscap quenches.” We then investigated the time evolution of entanglement in both general CFT and holographic settings of crosscap quenches, with the latter serving as an analytically tractable example of chaotic systems. Additionally, we numerically studied crosscap quenches in both a non-integrable critical spin chain and an integrable critical spin chain.
In the non-integrable systems studied in this paper, local subsystems, such as small single intervals, exhibit a volume law for entanglement and are expected to remain thermal throughout the dynamics. On the other hand, the entanglement entropy of antipodally located double intervals initially exhibits an area law, followed by linear growth, and eventually saturates to a volume law. This behavior illustrates how the initially well-organized thermal state becomes scrambled under time evolution and transitions into a state of few-body thermalization. Conversely, the integrable system deviates from the initially well-organized thermal equilibrium and exhibits oscillations in the time evolution of entanglement.

### v2

@@
Understanding the mechanisms by which complex correlations emerge through the dynamics of quantum many-body systems remains a fundamental challenge in modern physics. To address this, quench dynamics starting from nonthermal states have been extensively studied, leading to significant progress. In this paper, we propose a novel quench protocol, termed the “crosscap quench”, to investigate how highly structured thermal pure states relax into typical ones. We begin by analyzing conformal field theories (CFTs) and derive universal features in the time evolution of the entanglement entropy. Furthermore, leveraging the AdS/CFT correspondence, we study holographic CFTs, providing an analytically tractable example in chaotic CFTs. Finally, we validate these findings through numerical simulations in both nonintegrable and integrable quantum spin systems.
@@ Introduction
While |$\ket{\rm EAP(\beta)}$ serves as a thermal pure state, it is a highly atypical one. In fact, it can be easily distinguished from a typical state by the antipodal correlation functions, which are defined on a few degrees of freedom but are considered highly nonlocal in 1D systems. In this sense, the EAP state utilizes the antipodal entanglement structure to deceive local observers into perceiving thermal equilibrium and at the same time stays easy to simulate.
As a result, even though |$\ket{\rm EAP(\beta)}$ is already in thermal equilibrium, its correlation structure will get more scrambled under chaotic quenches. We will study the time evolution of the entanglement entropy after such quenches. In contrast to correlation functions, the entanglement entropy is defined independently of the choice of fields, making it a useful tool for highlighting a hierarchy between different notions of thermal equilibrium. In particular, EAP states exhibit a remarkable property in which a single interval is maximally entangled with its complement, whereas an antipodally located double interval is completely separable from its complement. We will investigate how this sharp entanglement structure evolves toward that of a typical state. We will also study time evolution of EAP states in integrable systems to elucidate the role of nonintegrability of the system.
@@
Note added: While this paper is under preparation, a very interesting paper [9] came out and also studied time evolution of the entanglement entropy from crosscap states but in different systems. The authors of [9] focused on quantum circuit models and integrable fermion chains while we focus on CFT, holography and spin chains in this paper. The integrable fermion chains investigated in [9] should be equivalent to our computations for the transverse-field Ising models in section 5 via the Jordan-Wigner transformation in some regimes.
@@ 2.1 Entanglement Structure of Crosscap States
The first situation we would like to consider is a single interval subsystem $A$ given by $0\leq x\leq l$. At the thermodynamic limit $\alpha/L\ll 1$, in the regime $l/L\ll 1$,
@@ 3.4 Time Evolution of EE after the Crosscap Quench
Here, |$\ket{0}_{j}$ and |$\ket{1}_{j}$ are the eigenvectors of $\sigma_{j}^{z}$ in the local Hilbert space associated with site $j$. For this choice of the EAP state, as shown in Ref. [5], the imaginary-time evolved EAP state |$\ket{{\rm EAP}(\beta)}\propto e^{-\frac{\beta}{4}H}\ket{\rm EAP}$ is a thermal pure state at inverse temperature $\beta$ because the Hamiltonian (4.1) is real in the spin basis. For simplicity, we set $\beta=0$ in the following discussion. In this case, the entanglement entropy for a single interval and an antipodally located double interval, each with volume $l$, are exactly $\min\{l,2L-l\}\log 2$ and 00, respectively, simplifying the comparison between the pre- and post-quench states. Then the time-evolved state at time $t$ is given by
@@
Figure 1 shows the dynamics of entanglement entropy for the single interval $A=\{1,2,\dots,l\}$ and the antipodally located double interval $A=\{1,2,\dots,l/2\}\cup\{N+1,N+2,\dots,N+l/2\}$. For $l\gg L$, the entanglement entropy exhibits an initial linear growth and saturates at a constant value, in well agreement with CFT predictions. Moreover, this saturated value coincides with the thermodynamic entropy at $\beta=0$, $l\log 2$. These results implies that the EAP state acquires a complex correlation structure and relaxes into a few-body thermal state.
Figure 1: Entanglement entropy $S_{A}$ of the time-evolved EAP state |$\ket{{\rm EAP}(t)}$ in the critical Heisenberg chain with next-nearest-neighbor interactions as a function of $t$ for total system size $242N=$ 24 and various subsystem size $l$, with solid lines representing the antipodally located double intervals $A=\{1,2,\dots,l/2\}\cup\{N+1,N+2,\dots,N+l/2\}$ and dashed lines represent the single intervals $A=\{1,2,\dots,l\}$.
As illustrated in Fig. 2, the time-evolved EAP state |$\ket{{\rm EAP}(t)}$ contains a volume-law entanglement with a slight deviation increasing as the subsystem size grows, and the deviation becomes most pronounced when the subsystem size is exactly half of the total system size. Then we plot the system-size dependence of the deviation in Fig. 2. It can be seen that the deviation is subextensive. This is consistent with the results obtained from the analysis of holographic CFTs. Furthermore, we find that the deviation scales as $N^{0}$ in the thermodynamic limit.
Figure 2: (2) Time-averaged entanglement entropy of the time-evolved EAP state |$\ket{{\rm EAP}(t)}$ in the critical Heisenberg chain with next-nearest-neighbor interactions, averaged over the time interval $t\in[8,32]$, as a function of the subsystem size $l$ for the single interval and the antipodally located double interval in a system of size $202N=$ 20. The solid line represents the entanglement entropy of a single interval in the initial EAP state, which exhibits maximal entanglement for any such interval. (2) Deviation from the maximum value at $l=N$. The solid line serves as a guide to the eye and is proportional to $N^{-1}$.
@@
Figure 4: (4) Time-averaged entanglement entropy of the time-evolved EAP state |$\ket{{\rm EAP}(t)}$ in the critical transverse-field Ising chain, averaged over the time interval $t\in[8,32]$, as a function of the subsystem size $l$ for the single interval and the antipodally located double interval in a system of size $202N=$ 20. The solid line represents the entanglement entropy of a single interval in the initial EAP state, which exhibits maximal entanglement for any such interval. (4) Deviation from the maximum value at $l=N$.
@@ Conclusion
In this paper, we have studied quantum quenches in 1D many-body systems starting from crosscap states and EAP states. While EAP states are originally defined in spin systems, we have argued that crosscap states serve as good analogues of them in CFT because they exhibit similar entanglement structures. This was demonstrated by studying the universal behavior of the entanglement Rényi entropy of CFT crosscap states. Based on this, we termed the quenches starting from crosscap states and EAP states as “crosscap quenches.” We then investigated the time evolution of entanglement in both general CFT and holographic settings of crosscap quenches, with the latter serving as an analytically tractable example of chaotic systems. Additionally, we numerically studied crosscap quenches in both a nonintegrable critical spin chain and an integrable critical spin chain.
In the nonintegrable systems studied in this paper, local subsystems, such as small single intervals, exhibit a volume law for entanglement and are expected to remain thermal throughout the dynamics. On the other hand, the entanglement entropy of antipodally located double intervals initially exhibits an area law, followed by linear growth, and eventually saturates to a volume law. This behavior illustrates how the initially well-organized thermal state becomes scrambled under time evolution and transitions into a state of few-body thermalization. Conversely, the integrable system deviates from the initially well-organized thermal equilibrium and exhibits oscillations in the time evolution of entanglement.